Q1-You decide to hedge your position in the stock and buy options at the fair market value , when strike prices of 60$ .

a) what is the value of the option premium to hedge your position

b) what is the value of your position at the end of the year of the stock trades at 75$ per share , taking into consideration all costs associated with buying the options
1
Expert's answer
2015-11-03T00:00:47-0500
a) what is the value of the option premium to hedge your position
To calculate the option premium using this formula:
Option premium = Strike price - Going market rate
Image that a going market rate is $10, then:
Option premium = 60-10 = $50
b) what is the value of your position at the end of the year of the stock trades at 75$ per share , taking into consideration all costs associated with buying the options.
Value of position at the end = 75 – 50 = $25